Determine the critical value for a left-tailed test of a population mean at the α = 0.05 level of significance based on a sample size of n = 35.

A. 1.691

B. -2.728

C. 1.690

D. -1.691
Answer #1

Solution :

Given that,

This is the left tailed test

sample size = n = 35

Degrees of freedom = df = n - 1 = 35 - 1 = 34

\alpha = 0.05

t,df = t0.05,34 = -1.691

critical value = -1.691

option D. is correct